Find the third, fourth, and fifth terms in the sequence if
A1=3, a2=6 and an= 3an-1-2an-2

Answer by robertb(5830)

a3 = a2 - a1 = 6-3 = 3,
a4 = a3 - a2 = 3-6 = -3,
a5 = a4 - a3 = -3 - 3 = -6.